OpenStudy (anonymous):

Help basic Algebra!! find x^12

OpenStudy (anonymous):

If \[x ^{x^{6}} = \sqrt{2}^{2^{\frac{ 1 }{ 2 }}}\] find \[x ^{12}\] ___________ A)\[2\sqrt{2}\] B) 8 C) 2 D)\[4\sqrt{2}\] E)16 I need some help from NASA :D

OpenStudy (koikkara):

@arturo_prado I'm not sure will this link help you.. http://www.algebra.com/algebra/homework/equations/Equations.faq.question.182224.html Wondering If you are satisfied with this answer, Please Click "Best Response" and close this Question. Thank You ! Keep in touch with Open Study.. Bye !

OpenStudy (anonymous):

a is the answer